# MEMOIR: Lifelong Model Editing with Minimal Overwrite and Informed Retention for LLMs

### v1

@@ 4 Experiments
In this section, we present a quantitative evaluation of MEMOIR across different benchmarks, LLMs, and numbers of edits, ranging from $1$ to $7000$. The best average performance metric is in bold.
@@ 4.2 Main Experimental Results
Table 1 presents the results on the ZsRE zsre dataset for the question answering task, evaluated across a range of total edits from $1$ to $1000$. ZsRE is a context-free Q&A benchmark designed to assess a model’s ability to store and retrieve specific factual knowledge. We observe that MEMOIR consistently achieves the most balanced performance across all evaluation settings. On LLaMA-3, MEMOIR maintains a reliability above 0.95 and a generalization score above 0.91, with its locality metric reaching 1.0 across different edit counts. In contrast, other parametric methods exhibit substantial degradation as the number of edits increases. For example, WISE drops from an average metric of 0.92 to 0.77 as the number of edits grows from 1 to 1000. A similar trend is observed with AlphaEdit, which declines from 0.96 to 0.72. Although GRACE, a non-parametric baseline, maintains high reliability, it suffers from poor generalization, leading to a significantly lower average metric overall compared to MEMOIR. In summary, MEMOIR offers the best trade-off between reliability, generalization and locality, achieving an average metric of 0.95 on LLaMA-3 with 1000 total edits—outperforming all prior methods by a substantial margin of 0.16. Similar results are observed with Mistral, where MEMOIR again delivers the highest average metric, demonstrating its robustness and effectiveness across different LLMs.
@@
LLMs are prone to hallucinations—generating fluent but factually inaccurate or misleading content—which poses a significant challenge for their deployment in domains that demand high factual precision. We validate here the effectiveness of MEMOIR on correcting hallucinations on the SelfCheckGPT selfcheckgpt dataset, which consists of Wikipedia-style biographies generated by GPT-3 gpt-3 . The dataset is annotated for factual hallucinations, with ground-truth replacements verified against Wikipedia. We note that we report only locality and reliability using perplexity (PPL) metric, following (wang2024wise, ), as no suitable generalization metric exists in this setting. As shown in Table 2, MEMOIR again achieves the best balanced performance, with its advantages becoming even more significant with a larger number of edits. Under the most challenging setting of $600$ edits, MEMOIR maintains a saturated locality score while delivering a perplexity metric, 57% and by 77% smaller than WISE, the second-best performing approach, respectively on LLaMA-3 and on Mistral. These results further confirm MEMOIR’s robustness in sustaining model quality during long editing sequences.
@@ 4.3 Scaling with longer sequences of edits
Table 4: Editing performance on LLaMA-3 with 5000 and 7000 edits.
Method 5000 7000
@@
MEMIT (memit, ) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
GRACE (grace, ) 1.00 0.31 1.00 0.77 1.00 0.31 1.00 0.77
WISE (wang2024wise, ) 0.59 0.57 1.00 0.72 0.59 0.57 1.00 0.72
@@
MEMOIR (Ours) 0.91 0.88 0.99 0.93 0.90 0.86 0.99 0.92
The ultimate goal of lifelong model editing algorithm is to continually deliver reliable performance with increasingly large number of edits without performance degradation on previously edited knowledge. Here we extend the total number of edits to 5000 and 7000 edits on LLaMA-3, almost exploiting all samples from ZsRE dataset. Table 4 presents the performance of MEMOIR using LLaMA-3 model and shows MEMOIR’s remarkable scalability to total number of edits beyond prior baselines. Averaged across all metrics, MEMOIR remains by far the best performing method on both 5000 and 7000 edits, with a margin of over $0.15$ compared to prior methods. A full performance trajectory from 1000 to 3000 edits, comparing MEMOIR against key baselines on both LLaMA-3 and Mistral in provided in Figure 6 in Appendix.
@@ 5 Analysis and Discussion
We provide additional analysis of MEMOIR ’s performance. Unless noted otherwise, experiments are based on the LLaMA-3 model with 1000 sequential edits on ZsRE.
@@
MEMOIR 0.95 0.91 1.00 0.95
MEMOIR w/o K.A. 0.94 0.86 0.28 0.69
We evaluate the effectiveness of inference-time conditional knowledge activation strategy introduced in Section 3.2. Specifically, we directly use the mask based on the input activations without matching step of Equation 4 that enables either the activation of only relevant memory knowledge or the activation of the memory module altogether. Table 5 shows that disabling conditional knowledge activation results in a noticeable drop in generalization performance, since the model is unable to precisely activate the parameters that store relevant knowledge. Moreover, the absence of conditional activation significantly degrades locality: performance on irrelevant prompts drops below 0.3, whereas MEMOIR maintains full performance on these samples by relying solely on pre-trained knowledge.
@@
MEMOIR distributes knowledge updates across distinct parameter subsets to reduce interference and mitigate forgetting of previously edited samples. To quantify this effect, we report the post-edit accuracy for each sample in the editing sequence. As shown in Figure 6, MEMOIR exhibits the smallest degradation in accuracy for earlier edits. Specifically, it maintains an average accuracy of approximately 0.92 for the first 100 edits, even after all 1000 edits have been applied. In contrast, the accuracy of the second-best method, AlphaEdit, on the first 100 edits drops to 0.71. This pattern holds across the full sequence, highlighting MEMOIR’s ability to preserve earlier edits, while baseline methods experience significant forgetting.
@@ A.1 Datasets
Irrelevant prompt and target $\bm{x}_{\text{e}},\bm{y}_{ood}$ Adam Silver, the current commissioner of the National Basketball Association (NBA), has made a significant impact on the league since taking the position in 2014. Prior to becoming the commissioner, Silver had already established a solid foundation within the NBA, holding various roles including chief operating officer and deputy commissioner. Under his leadership, the NBA has experienced considerable growth, both economically and globally. One particular milestone during Silver’s tenure was his handling of the Donald Sterling controversy in 2014, where he forced Sterling to sell the Los Angeles Clippers and banned him from all NBA games and events. Silver’s decisive action showcased his commitment to maintaining the integrity of the NBA and addressing issues of racism within the league.
@@ A.2 Metrics
The metrics for Q&A and hallucination correction are introduced in the main paper Sec. 4.1. Here, we detail the implementation specific to the Temporal dataset. We adopt the standard evaluation protocol based on reliability on Temporal dataset as performed for both Q&A and hallucination correction tasks.
@@
[…]
@@ A.4.1 Training Details
All experiments are conducted on a single NVIDIA A100 GPU. Editing 1,000 Q&A samples with MEMOIR takes approximately 3 hours for LLaMA-3 and 4 hours for Mistral, using an SGD optimizer. For the hallucination dataset, editing 600 samples requires around 4 hours on LLaMA-3 and 6 hours on Mistral.
@@ A.4.2 Baseline implementations
For ROME, we follow the default setting in their sourcecode on GPT-J, as well as the implementation in [14] and [9], to edit the 5th layer in the LLM for both LLaMA-3 and Mistral. The covariance statistics are computed using 100,000 samples from Wikitext in fp32 precision.
@@
For both LLaMA and Mistral models, MEMIT updates layers [4, 5, 6, 7, 8] and sets $\lambda$, the covariance adjustment factor, to 15,000, following the settings in [9]. Following ROME [11], it uses 100,000 Wikipedia samples.
@@
For AlphaEdit, we adopt the same hyperparameter settings as in [15] to ensure fair comparison. Specifically, for the LLaMA-3 (8B) model, we edit critical layers [4, 5, 6, 7, 8]. To compute hidden representations at each critical layer, we perform 25 optimization steps with a learning rate of 0.1. For LLaMA-2 and Mistral, we apply the same configuration. For the GPT-J model, we target layers [3, 4, 5, 6, 7, 8], and similarly perform 25 optimization steps per layer with a learning rate of 0.5. As in MEMIT [14], it uses $\lambda=15,000$ and 100,000 Wikipedia samples.
@@ A.4.3 Implementation details for MEMOIR
For all model architectures, including LLaMA-2, LLaMA-3, and Mistral, we edit the layer model.layers[27].mlp.down_proj.weight. We use the SGD optimizer with a learning rate of 1.0 and a gradient clipping with clipping threshold of 1.0. The number of iterations per edit is set as 70 for LLaMA-2, and Mistral, and reduces to 30 for LLaMA-3 for accelerating computation.
The number of active indices $k$ for TopHash is set as 4096 for LLaMA-3, LLaMA-2, and Mistral architecture. The threshold for conditional knowledge activation $\tau$ is set as 0.6 for LLaMA-3 and LLaMA-2, and 0.55 for Mistral.
@@ A.4.4 Experiments for a small number of edits
In Table 1 and Table 2, the total number of edits gradually increases from $1$ to $1000$ for the ZsRE dataset and from $1$ to $600$ for the SelfcheckGPT dataset. To reduce the variance introduced by individual edit samples in experiments involving a small number of edits ($T=1$ to $T=100$), we report results averaged over multiple runs using different edit samples. For example, results for $T=1$ on the ZsRE dataset are averaged over 1,000 independent experiments, each using a different edit sample. Similarly, for $T=10$ on the SelfcheckGPT dataset, we average results over 60 runs with distinct edit samples. In total, all experiments span 1,000 edit samples on ZsRE and 600 on SelfcheckGPT.
@@ B.1 Model editing on LLaMA-2-7B
We provide in this section the result for editing LLaMA-2-7B model with MEMOIR on both Q&A and hallcination correction tasks. We demonstrate that MEMOIR continues to deliver state-of-the-art performance, similarly to the results on LLaMA-3 and Mistral in main text.
@@
MEMOIR (Ours) 1.00 0.92 1.00 0.97 0.99 0.93 1.00 0.97 0.98 0.91 1.00 0.96 0.95 0.89 1.00 0.95
We present the results on the ZsRE dataset with LLaMA-2 models in Table 9. The results demonstrate that MEMOIR delivers the best performance against the compared baseline approaches, where its advantage becomes increasingly pronounced with larger numbers of total edits. For a total of 1000 edits, MEMOIR achieves an average score of $0.95$ across the three evaluation metrics, namely reliability, generalization, and locality, outperforming the second-best method, WISE, by a margin of $0.12$. Notably, while most baselines show significant performance degradation as the number of edits increases—particularly in generalization and locality—MEMOIR maintains robust performance across all three axes.
@@
MEMOIR (Ours) 1.00 1.00 1.01 1.00 1.03 1.00 1.09 1.00
We report results on the Hallucination setting using the SelfCheckGPT dataset with LLaMA-2 in Table 10. Across all edit scales, MEMOIR consistently outperforms baseline methods in both reliability (lower perplexity) and locality (higher consistency with the original model). Notably, while several baselines suffer significant degradation in reliability as the number of edits increases, e.g., ROME and MEMIT reaching over 100 in perplexity at $T=600$, MEMOIR maintains low perplexity (1.09) and perfect locality (1.00). This indicates that our method successfully incorporates new knowledge without disrupting unrelated predictions. On average across all edit scales and metrics, MEMOIR demonstrates a robust scalability in correcting hallucinations.
B.2 Additional ablations

### v2

@@ 4 Experiments
In this section, we present a quantitative evaluation of MEMOIR across different benchmarks, LLMs, and numbers of edits, ranging from $1$ to 15,00015{,}000. The best average performance metric is in bold.
@@ 4.2 Main Experimental Results
Table 1 presents the results on the ZsRE zsre dataset for the question answering task, evaluated across a range of total edits from $1$ to $1000$. ZsRE is a context-free Q&A benchmark designed to assess a model’s ability to store and retrieve specific factual knowledge. We observe that MEMOIR consistently achieves the most balanced performance across all evaluation settings. On LLaMA-3, MEMOIR maintains a reliability of 0.94 and a generalization of 0.85, with its locality reaching 1.0 across different edit counts. In contrast, other parametric methods exhibit substantial degradation as the number of edits increases. For example, WISE drops from an average metric of 0.92 to 0.77 as the number of edits grows from $1$ to 1,0001{,}000. A similar trend is observed with AlphaEdit, which declines from 0.96 to 0.72. Although GRACE, a non-parametric baseline, maintains high reliability, it suffers from poor generalization, leading to a significantly lower average metric overall compared to MEMOIR. In summary, MEMOIR offers the best trade-off between reliability, generalization, and locality, achieving an average metric of 0.93 on LLaMA-3 with 1,0001{,}000 total edits—outperforming all prior methods by a substantial margin of 0.14. Similar results are observed with Mistral, where MEMOIR again delivers the highest average metric, demonstrating its robustness and effectiveness across different LLMs.
@@
LLMs are prone to hallucinations—generating fluent but factually inaccurate or misleading content—which poses a significant challenge for their deployment in domains that demand high factual precision. We validate here the effectiveness of MEMOIR on correcting hallucinations on the SelfCheckGPT selfcheckgpt dataset, which consists of Wikipedia-style biographies generated by GPT-3 gpt-3 . The dataset is annotated for factual hallucinations, with ground-truth replacements verified against Wikipedia. We note that we report only locality and reliability using perplexity (PPL) metric, following (wang2024wise, ), as no suitable metric for generalization exists in this setting. As shown in Table 2, MEMOIR again achieves the best balanced performance, with its advantages becoming even more significant with a larger number of edits. Under the most challenging setting of $600$ edits, MEMOIR maintains a saturated locality while delivering a perplexity metric, 61% and by 77% smaller than WISE, the second-best performing approach, respectively on LLaMA-3 and on Mistral. These results further confirm MEMOIR’s robustness in sustaining model quality during long editing sequences.
@@
Multi-hop reasoning
In appendix B.5 we assess the efficacy of MEMOIR in multi-hop reasoning, a more challenging scenario. We evaluate on the RippleEdits benchmark (ripple), where MEMOIR demonstrates better multi-hop reasoning generalization than all baseline methods.
@@ 4.3 Scaling with longer sequences of edits
Table 4: Editing performance on LLaMA-3 with 7,0007{,}000 and 15,00015{,}000 edits.
Method T=7, [^0] T={7{,}000} T=15, [^0] T={15{,}000}
@@
MEMIT (memit, ) 0.00 0.00 0.00 0.00
GRACE (grace, ) 1.00 0.31 1.00 0.77 0.99 0.28 1.00 0.76
WISE (wang2024wise, ) 0.59 0.57 1.00 0.72 0.43 0.41 1.00 0.61
@@
MEMOIR (Ours) 0.92 0.85 0.99 0.92 0.87 0.78 0.98 0.88
The ultimate goal of lifelong model editing algorithm is to continually deliver reliable performance with an increasingly large number of edits without performance degradation on previously edited knowledge. Here we extend the total number of edits to 7,0007{,}000 and 15,00015{,}000 edits on LLaMA-3, almost exploiting all samples from ZsRE dataset. Table 4 presents the performance of MEMOIR using LLaMA-3 model and shows MEMOIR’s remarkable scalability to the total number of edits beyond prior baselines. Averaged across all metrics, MEMOIR remains by far the best performing method on both 7,0007{,}000 and 15,00015{,}000 edits, with a margin of over 0.12 compared to prior methods. A full performance trajectory from 1,0001{,}000 to 3,0003{,}000 edits, comparing MEMOIR against key baselines on both LLaMA-3 and Mistral is provided in Figure 6 in Appendix.
@@ 5 Analysis and Discussion
We provide additional analysis of MEMOIR ’s performance. Unless noted otherwise, experiments are based on the LLaMA-3 model with 1,0001{,}000 sequential edits on ZsRE.
@@
We evaluate the effectiveness of inference-time conditional knowledge activation strategy introduced in Section 3.2. Specifically, we directly use the mask based on the input activations without the matching step of Equation 4 that enables either the activation of only relevant memory knowledge or the activation of the memory module altogether.
@@
MEMOIR 0.94 0.85 1.00 0.93
MEMOIR w/o K.A. 0.94 0.77 0.68 0.80
Table 5 shows that disabling conditional knowledge activation results in a noticeable drop in generalization performance, since the model is unable to precisely activate the parameters that store relevant knowledge. Moreover, the absence of conditional activation significantly degrades locality: performance on irrelevant prompts drops below 0.7, whereas MEMOIR maintains full performance on these samples by relying solely on pre-trained knowledge.
@@
MEMOIR distributes knowledge updates across distinct parameter subsets to reduce interference and mitigate forgetting of previously edited samples. To quantify this effect, we report the post-edit accuracy for each sample in the editing sequence. As shown in Figure 6, MEMOIR exhibits the smallest degradation in accuracy for earlier edits. Specifically, it maintains an average accuracy of near 0.90 for the first 100 edits, even after all $1000$ edits have been applied. In contrast, the accuracy of the second-best method, AlphaEdit, on the first 100 edits drops to 0.71. This pattern holds across the full sequence, highlighting MEMOIR’s ability to preserve earlier edits, while baseline methods experience significant forgetting.
@@ A.1 Datasets
OOD prompt and target $\bm{x}_{\text{e}},\bm{y}_{\text{ood}}$ Adam Silver, the current commissioner of the National Basketball Association (NBA), has made a significant impact on the league since taking the position in 2014. Prior to becoming the commissioner, Silver had already established a solid foundation within the NBA, holding various roles including chief operating officer and deputy commissioner. Under his leadership, the NBA has experienced considerable growth, both economically and globally. One particular milestone during Silver’s tenure was his handling of the Donald Sterling controversy in 2014, where he forced Sterling to sell the Los Angeles Clippers and banned him from all NBA games and events. Silver’s decisive action showcased his commitment to maintaining the integrity of the NBA and addressing issues of racism within the league.
Irrelevant prompt $\bm{x}_{\text{irr}}$ Written by Gerald H. Nelson 2̆022 Fred B. Burch Album This song officially appears on the Red Rose Speedway (2018) Official album. Sessions This song has
Irrelevant target $\bm{y}_{\text{irr}}$ been recorded during the following sessions Red Rose Speedway sessions at Olympic Studios (March 1972)Other unreleased songs from the double-album Re
@@ A.2 Metrics
The metrics for Q&A and hallucination correction are introduced in the main paper Sec.˜4.1. Here, we detail the implementation specific to the Temporal dataset. We adopt the standard evaluation protocol based on reliability on Temporal dataset as performed for both Q&A and hallucination correction tasks.
@@
$\text{OOD Gen.}=\frac{1}{T}{\sum_{t=1}^{T}}\mathbbm{1}\{L_{\bm{\theta}}^{T}(\bm{x}_{\text{e}}^{t},\bm{y}_{\text{ood}}^{t})<\delta\},$
@@ A.4.1 Training Details
All experiments are conducted on a single NVIDIA A100 GPU. Editing 1,0001{,}000 Q&A samples with MEMOIR takes approximately 3 hours for LLaMA-3 and 4 hours for Mistral, using an SGD optimizer. For the hallucination dataset, editing 600 samples requires around 4 hours on LLaMA-3 and 6 hours on Mistral.
@@ A.4.2 Baseline implementations
For ROME, we follow the default setting in their sourcecode on GPT-J, as well as the implementation in [14] and [9], to edit the 5th layer in the LLM for both LLaMA-3 and Mistral. The covariance statistics are computed using 100,000100{,}000 samples from Wikitext in fp32 precision.
@@
For both LLaMA and Mistral models, MEMIT updates layers [4, 5, 6, 7, 8] and sets $\lambda$, the covariance adjustment factor, to 15,00015{,}000, following the settings in [9]. Following ROME [11], it uses 100,000100{,}000 Wikipedia samples.
@@
For AlphaEdit, we adopt the same hyperparameter settings as in [15] to ensure fair comparison. Specifically, for the LLaMA-3 (8B) model, we edit critical layers [4, 5, 6, 7, 8]. To compute hidden representations at each critical layer, we perform 25 optimization steps with a learning rate of 0.1. For LLaMA-2 and Mistral, we apply the same configuration. For the GPT-J model, we target layers [3, 4, 5, 6, 7, 8], and similarly perform 25 optimization steps per layer with a learning rate of 0.5. As in MEMIT [14], it uses $\lambda=15{,}000$ and 100,000100{,}000 Wikipedia samples.
@@ A.4.3 Implementation details for MEMOIR
For all model architectures, including LLaMA-2, LLaMA-3, Mistral, and GPT-J, we edit the layer model.layers[27].mlp.down_proj.weight for LLaMA-2, LLaMA-3 and Mistral, and transformer.h[21].mlp.fc_out.weight for GPT-J. We use the SGD optimizer with a learning rate of 1.0 and a gradient clipping with clipping threshold of 1.0. The number of iterations per edit is set as 70 for LLaMA-2, Mistral, and GPT-J, and reduces to 30 for LLaMA-3 for accelerating computation.
The number of active indices $k$ for TopHash is set as 4096 for LLaMA-3, LLaMA-2, Mistral, and GPT-J architecture. The threshold for conditional knowledge activation $\tau$ is set as 0.4 for LLaMA-3 and Mistral, 0.46 for LLaMA-2, and 0.45 for GPT-J.
@@ A.4.4 Experiments for a small number of edits
In Table 1 and Table 2, the total number of edits gradually increases from $1$ to 1,0001{,}000 for the ZsRE dataset and from $1$ to $600$ for the SelfcheckGPT dataset. To reduce the variance introduced by individual edit samples in experiments involving a small number of edits ($T=1$ to $T=100$), we report results averaged over multiple runs using different edit samples. For example, results for $T=1$ on the ZsRE dataset are averaged over 1,0001{,}000 independent experiments, each using a different edit sample. Similarly, for $T=10$ on the SelfcheckGPT dataset, we average results over 60 runs with distinct edit samples. In total, all experiments span 1,0001{,}000 edit samples on ZsRE and 600 on SelfcheckGPT.
@@ B.1 Model editing on LLaMA-2-7B
We provide in this section the result for editing LLaMA-2-7B model with MEMOIR on both Q&A and hallucination correction tasks. We demonstrate that MEMOIR continues to deliver state-of-the-art performance, similarly to the results on LLaMA-3 and Mistral in main text.
@@
MEMOIR (Ours) 1.00 0.96 1.00 0.99 0.99 0.92 1.00 0.97 0.98 0.91 1.00 0.96 0.97 0.90 1.00 0.96
We present the results on the ZsRE dataset with LLaMA-2 models in Table 9. The results demonstrate that MEMOIR delivers the best performance against the compared baseline approaches, where its advantage becomes increasingly pronounced with larger numbers of total edits. For a total of 1,0001{,}000 edits, MEMOIR achieves an average score of $0.95$ across the three evaluation metrics, namely reliability, generalization, and locality, outperforming the second-best method, WISE, by a margin of $0.13$. Notably, while most baselines show significant performance degradation as the number of edits increases—particularly in generalization and locality—MEMOIR maintains robust performance across all three axes.
@@
MEMOIR (Ours) 1.16 1.00 1.09 1.00 1.17 1.00 1.48 0.99
We report results on the Hallucination setting using the SelfCheckGPT dataset with LLaMA-2 in Table 10. Notably, while several baselines suffer significant degradation in reliability as the number of edits increases, e.g., ROME and MEMIT reaching over 100 in perplexity at $T=600$, MEMOIR maintains low perplexity (1.48) and high locality =(0.99). This indicates that our method successfully incorporates new knowledge without disrupting unrelated predictions. On average across all edit scales and metrics, MEMOIR demonstrates a robust scalability in correcting hallucinations.
B.2 Activation centering during mask construction
As described in the main text, the mask selection procedure involves a centering step, where a reference activation vector is subtracted from the activations of the input prompt $\bm{x}$. Specifically, this centering step suppresses features that exhibit consistently high magnitudes across different prompts, allowing the mask to focus on more distinctive features for each prompt. In this section, we describe the computation of the reference vector.
Table 11: Effect of varying the number of irrelevant prompts for activation centering on three post-edit metrics. Results are for experiments with 1,0001{,}000 edits on ZsRE dataset using LLaMA-3 model.
 Number of Irrelevant Prompts Rel. Gen. Loc. Avg.
0 (no centering) 0.92 0.85 1.00 0.92
$10$ 0.93 0.85 1.00 0.93
$100$ 0.94 0.85 1.00 0.93
1,0001{,}000 0.93 0.85 1.00 0.93
The irrelevant prompts used to compute the reference mean are drawn from the ZsRE dataset, ensuring no overlap with the edited prompts. We perform an ablation study by varying the number of irrelevant prompts used for mean estimation, ranging from 10 to 1,0001{,}000. We also include the results for not performing such feature centering process.
Results in Table 11 demonstrate that the performance of MEMOIR is robust to the number of irrelevant prompts used for decentering. We further note that even without the decentering step, MEMOIR still demonstrates state-of-the-art performance with a large performance margin to previous methods.
B.3 Additional ablations